OpenStudy (anonymous):

The second angle of a triangle is 3 times as large as the first. The third angle is more than the first. Find all three angles.

OpenStudy (anonymous):

Welcome back. Ready to work together?

OpenStudy (anonymous):

of course

myininaya (myininaya):

First angle=F Second angle=S Third angle=T F+S+T=180 S=3*F T>F

myininaya (myininaya):

i think you are missing something

OpenStudy (anonymous):

Oh! Looks like myininaya's got this. He/she is good at this stuff. I'll pass it to them. :P

myininaya (myininaya):

the third is more than the first?

OpenStudy (anonymous):

no that is how the question was asked

OpenStudy (anonymous):

It's missing information then. :(

OpenStudy (anonymous):

so it cant be solved

myininaya (myininaya):

well we have F+S+T=180 and S=3*F F+3*F+T=180 <- we have two unknowns in this equation

OpenStudy (anonymous):

you would need to know how much more than the first the third angle is.

myininaya (myininaya):

4F+T=180

myininaya (myininaya):

the third angle is ____ more than the first

OpenStudy (anonymous):

how do i figure that out

myininaya (myininaya):

is there something that should be in that blank

OpenStudy (anonymous):

its not something you should figure out, it should be given to you. otherwise there isnt enough info.

OpenStudy (anonymous):

its a question that was asked in my final

OpenStudy (anonymous):

(It's like saying "Bill is 2 years older than Bob. How old is Bob?" Unless they tell you how old Bill is, we can't know how old Bob is.)